OpenStudy (anonymous):

Lara chooses a square number. She rounds it to the nearest hundred. Her answer is 200 Write all the possible square numbers Lara could have chosen.

OpenStudy (anonymous):

12*12=144 13*13=169 is 200 rounded to the nearest hundred 14*14=196 is 200 rounded to the nearest hundred 15*15=225 is 200 rounded to the nearest hundred 16*16=256 is 300 rounded to the nearest hundred 17*17=289 So, take your choice.....